At 373 K, Kp = 0.416 for the equilibrium 2 NOBr(g) -> 2NO(g) + Br2(g). If the pressures of NOBr(g) and NO(g) are equal, what is the equilibrium pressure of Br2(g)?
1
Expert's answer
2018-03-06T07:53:15-0500
2 NOBr(g) -> 2NO(g) + Br2(g).
If p(Br2) = x, than p(NOBr) = p(NO) = 2x
Kp = p(Br2)2* p(NO)/ p(NOBr)2
Kp = (2x)2*x/(2x)2 = x = 0.416 Pa